# Emergent Hydrodynamics in an Exclusion Process with Long-Range Interactions

### v1

@@ 1 Introduction
Here $j^{\ast}(\cdot)$ is the stationary current density relation which like the collective diffusion coefficient $D(\cdot)$ depend only on the local density $\rho(x,t)$ [1, 2]. Thus the hydrodynamic equation (2) is a closed (in general nonlinear) deterministic partial differential equation for the macroscopic density profile $\rho(x,t)$. The macroscopic space point $x$ represents a rescaled microscopically large region of the underlying lattice on which the microscopic model is defined. More precisely, this rescaling amounts to introducing a lattice constant $a$ and looking at the particle system in a region on the lattice of length $na$ around some site. In the limit $a\to 0$ this lattice site and the region around it then represent the space point $x$. The macroscopic time is also large compared to the microscopic time and usually taken with a rescaling $t\rightarrow t/a^{x}$ for some suitably chosen scaling exponent $x$ that depends on the particle system under consideration.
@@
Local stationarity also implies that the microscopic details of the initial distribution are washed out and do not matter. Therefore one expects that only the coarse-grained initial density $\rho_{0}(x):=\rho(x,0)$ determines the coarse-grained density $\rho(x,t)$ at a later macroscopic time $t$. Since the two properties of the law of large numbers in a large region and local stationarity are generic, the hydrodynamic limit (2) is expected to be valid generically, even when there is no rigorous mathematical proof [1]. Indeed, even when – like in boundary-driven systems – correlations are not short-ranged but sufficiently weak, the hydrodynamic description is still valid and can proved rigorously for some classes of stochastic interacting particle systems [5, 6, 7].
@@ 2.1 Definition of the SDEP
In the SDEP the interaction is long-ranged. The $i$-th particle at position $k_{i}$ attempts to jump to the neighboring site $k_{i}\pm 1$ with the rate
@@
which depends on the position of all other particles on the lattice. The factor $w$ sets the time scale of the process. The corresponding instantaneous current, i.e., the net number of particle jumps across the lattice bond $(k,k+1)$ in an infinitesimal time interval, is given by
@@
and where $Z_{N}$ is a normalization factor [15]. The above-mentioned link with the eigenvalues $\exp{(2\pi ix_{j})}$ of random unitary matrices that perform Dyson’s Brownian motion over the unitary group $U(N)$ becomes apparent by the identification $k_{j}:=x_{j}L$ of the parameters $x_{j}$ with the rescaled particle coordinates $k_{j}$. However we stress that here the eigenvalues are restricted to the discrete set of $L^{\rm th}$ roots of unity.
@@ 2.2 Some useful facts about the spin-1/2 XX quantum chain
$\Psi_{N}(\boldsymbol{k})=\frac{2^{N\choose 2}}{L^{N/2}}\prod_{i=1}^{j}\prod_{j=i+1}^{N}\sin{\left(\pi\frac{k_{j}-k_{i}}{L}\right)}.$ (14)
The corresponding ground state energy is [30]
@@ 3 Hydrodynamics
Note that, in the conventional SSEP, the spectral gap of the intensity matrix scales as $O(1/L^{2})$ corresponding to diffusive scaling with dynamical exponent $z=2$. This is of course consistent with the fact that the hydrodynamic limit of the SSEP is given by the heat equation [3]. In contrast, the spectral gap of the intensity matrix of the SDEP scales as $O(1/L)$, see Eq. (16), which leads to faster relaxation with dynamical exponent $z=1$. We thus expect that, upon coarse-graining under ballistic (Eulerian) scaling the SDEP satisfies a hydrodynamic equation resulting from the continuity equation
@@
3.1 Motivation for the conjectured particle current (25)
Let us start from the exact expression for the instantaneous current in the SDEP, Eq. (5). For a given configuration of particles at positions $k_{1},\dots,k_{N}$, we define the instantaneous density of particles on site $k$ as $n^{\rm inst}_{k}=\sum_{j=1}^{N}\delta_{k,k_{j}}\,\in\,\{0,1\}$. We rewrite formula (5) as
$j^{\rm inst}_{k}$ $=$ $\frac{w}{2}\,n^{\rm inst}_{k}(1-n^{\rm inst}_{k+1})\exp\left[\sum_{q\neq k,k+1}n^{\rm inst}_{q}\log\left(\frac{\sin{\pi\frac{q-k-1}{L}}}{\sin{\pi\frac{q-k}{L}}}\right)\right]$ (27)
$\qquad-\frac{w}{2}\,(1-n^{\rm inst}_{k})n^{\rm inst}_{k+1}\exp\left[-\sum_{q\neq k,k+1}n^{\rm inst}_{q}\log\left(\frac{\sin{\pi\frac{q-k-1}{L}}}{\sin{\pi\frac{q-k}{L}}}\right)\right].\qquad$
This expression is exact. Now we start making approximations. First, we can approximate the sum in the two exponentials in terms of the coarse-grained density $\rho(x)$,
$\sum_{\underset{q\neq k,k+1}{q=0}}^{L-1}n^{\rm inst}_{q}\log\left(\frac{\sin{\pi\frac{q-k-1}{L}}}{\sin{\pi\frac{q-k}{L}}}\right)$ $\underset{L\gg 1}{\simeq}$ $\left(\int_{0}^{k-\epsilon}+\int_{k+\epsilon}^{L}\right)\log\left(\frac{\sin{\pi\frac{x-k-1}{L}}}{\sin{\pi\frac{x-k}{L}}}\right)\rho(x)dx$
$\simeq$ ${\rm p.v.}\int_{0}^{L}\log\left(1-\frac{\pi}{L}{\rm cotan}\,\pi\frac{x-k}{L}\right)\rho(x)dx$
$\simeq$ $\pi\,\mathcal{H}\rho(k)\,.$
When we make this approximation, we expect that we are correctly describing the contribution of the long-distance terms with $|q-k|\gg 1$. However we are overlooking details at short distances, where $q$ is distant from $k$ from only a few lattice sites.
This suggests a second approximation: that the ratio
$\frac{w}{2}\frac{\,n^{\rm inst}_{k}(1-n^{\rm inst}_{k+1})\exp\left[\sum_{q\neq k,k+1}n^{\rm inst}_{q}\log\left(\frac{\sin{\pi\frac{q-k-1}{L}}}{\sin{\pi\frac{q-k}{L}}}\right)\right]}{\exp(\pi\mathcal{H}\rho(k))}$
depends only on the local state of the gas. This is a mean-field approximation. Upon coarse-graining, that ratio should become a function of the local particle density $\rho(x,t)$ only. Thus, a natural Ansatz for the coarse-grained particle current in the SDEP is
@@ 3.1 Motivation for the conjectured particle current (25)
for some functions $a_{+}(\cdot)$ and $a_{-}(\cdot)$.
Because of reflexion symmetry, we must have $a_{+}(\cdot)=a_{-}(\cdot)$. Moreover, if we look back at formula (27) and replace the minus sign between the two terms by a plus sign, we get the instantaneous local activity $a^{\rm inst}_{k}$ instead of the particle current $j^{\rm inst}_{k}$. So the coarse-grained activity should be
$a[\rho](x,t)\,=\,a_{+}(\rho(x,t))e^{\pi\mathcal{H}\rho(x,t)}\,+\,a_{-}(\rho(x,t))e^{-\pi\mathcal{H}\rho(x,t)}.$ (29)
Now we appeal to local equilibrium. When we apply this formula to an equilibrium state, which has a constant density $\rho(x,t)=\rho$, we have $\mathcal{H}\rho(x,t)=0$, and so we find that $a_{+}(\rho)+a_{-}(\rho)$ is the activity at equilibrium. Thus, we conclude that the coarse-grained activity and the coarse-grained current should be given respectively by
$\begin{array}[]{rcl}a[\rho](x,t)&=&a_{{\rm eq}}(\rho(x,t))\,\cosh(\pi\mathcal{H}\rho(x,t)),\\ \\ j[\rho](x,t)&=&a_{{\rm eq}}(\rho(x,t))\,\sinh(\pi\mathcal{H}\rho(x,t)),\end{array}$ (30)
where $a_{{\rm eq}}(\rho)$ is the local activity at equilibrium. The local activity at equilibrium should be identified with the exact lattice result (23), which leads to the conjectured form of the current (25).
@@ 3.2 Equivalence between non-local one-component hydrodynamics and local two-component hydrodynamics
as a local two-component hydrodynamic equation, with two densities $\rho(x)$ and $\tilde{\rho}(x)$ that can be viewed as independent,
@@
The reason for this equivalence is as follows (see also the appendix of Ref. [31] for a similar calculation). The fundamental property of the Hilbert transform is that $f(x)$ and $\mathcal{H}f(x)$ are the real- and imaginary parts of the same analytic function in the upper half-plane. More specifically, for a periodic function $f(x)$ of period $L$, there is exists a function $F(z)$ defined for any ${\rm Im}\,z>0$, that is analytic in the upper half-plane (including at infinity), with the same periodicity as $f$, i.e. $F(z+L)=F(z)$, and such that for all $x\in\mathbb{R}$
@@ 3.3 Relation with Abanov’s lattice free fermion imaginary-time hydrodynamics
The complex partial differential equation (35) is not new. In the context of large-scale descriptions of systems of hard-core particles or spin chains in one dimension, it was discussed by Abanov in Ref. [32]. In particular as pointed out by Abanov, for the quantity $\zeta(z,t)=\sin P(z,t)$ , it is equivalent to a complex Hopf (or inviscid Burgers) equation:
@@
The real-time hydrodynamics of the quantum XX spin chain can be easily understood, as it reflects the dynamics of the underlying non-interacting fermions [33]. One way to proceed is to think of the non-interacting fermions semi-classically. One can describe the system on large scales by its phase-space occupation $W(x,p)$, where the momentum $p$ is defined modulo $2\pi$ because the model is considered on the lattice.
Semi-classically, $W(x,p)$ is a function that takes values between 0 and $\frac{1}{2\pi}$. The local particle density is obtained by integrating over momenta at a fixed position $x$,
@@ 4.1 Density Profiles
Here $\ln$’ stands for the principal logarithm, with the branch cut along the negative real axis, i.e. $\ln z=\ln|z|+i\,\arg z$.
@@ 4.2 Limit shapes in spacetime
In Fig. 1 we see that the regions where the density is initially 0 or $1$ do not immediately melt. Instead these two regions have sharp boundaries that persists up to some time. Thus, when we look at the trajectories of particles in space-time for the SDEP, one clearly sees the appearance of a limit shape phenomenon, see Fig. 2. We observe the emergence of a sharp boundary between ‘frozen’ regions where the particle density is exactly 0 or $1$, and ‘fluctuating’ regions where the density is strictly between 0 and $1$. Such limit shape phenomena occur in various statistical physics problems, including crystal growth [53, 54], dimer coverings [55, 56, 57], statistics of random Young tableaux [58], the six-vertex model with domain-wall boundary conditions [59, 60, 61, 40], see e.g. [62, 63] for introductions. The curve separating the frozen regions and fluctuating regions is known as the arctic curve [55, 64, 65, 66, 67]. The hydrodynamic equation (31) can be exploited to make a prediction for the arctic curve for the above initial conditions, which can then be compared to the stochastic simulation of the SDEP, providing further numerical evidence for the validity of the hydrodynamic equation (31). This is what we do next.

### v2

@@ 1 Introduction
Here $j^{\ast}(\cdot)$ is the stationary current density relation which like the collective diffusion coefficient $D(\cdot)$ depend only on the local density $\rho(x,t)$ [72, 47]. Thus the hydrodynamic equation (2) is a closed (in general nonlinear) deterministic partial differential equation for the macroscopic density profile $\rho(x,t)$. The macroscopic space point $x$ represents a rescaled microscopically large region of the underlying lattice on which the microscopic model is defined. More precisely, this rescaling amounts to introducing a lattice constant $a$ and looking at the particle system in a region on the lattice of length $na$ around some site. In the limit $a\to 0$ this lattice site and the region around it then represent the space point $x$. The macroscopic time is also large compared to the microscopic time and usually taken with a rescaling $t\rightarrow t/a^{z}$ for some suitably chosen scaling exponent $z$ that depends on the particle system under consideration.
@@
Local stationarity also implies that the microscopic details of the initial distribution are washed out and do not matter. Therefore one expects that only the coarse-grained initial density $\rho_{0}(x):=\rho(x,0)$ determines the coarse-grained density $\rho(x,t)$ at a later macroscopic time $t$. Since the two properties of the law of large numbers in a large region and local stationarity are generic, the hydrodynamic limit (2) is expected to be valid generically, even when there is no rigorous mathematical proof [72]. Indeed, even when – like in boundary-driven systems – correlations are not short-ranged but sufficiently weak, the hydrodynamic description is still valid and can be proved rigorously for some classes of stochastic interacting particle systems [28, 9, 34].
@@ 2.1 Definition of the SDEP
In the SDEP the interaction is long-ranged. The $i$-th particle at is the average position $k_{i}$ attempts to jump to the neighboring site $k_{i}\pm 1$ with the rate
@@
which depends on the position of all other particles on the lattice. The factor $w$ sets the time scale of the process. For a fixed particle configuration $\boldsymbol{k}=(k_{1},\dots,k_{N})$, the quantity $j_{k}^{\rm inst}(\boldsymbol{k})$ denotes the microscopic instantaneous current across the bond $(k,k+1)$, i.e. the net jump rate from $k$ to $k+1$ minus the net jump rate from $k+1$ to $k$. It is given by
@@
and where $Z_{N}$ is a normalization factor [65]. Equation (4) is the natural nearest-neighbor reversible dynamics associated with the potential (7). Indeed, if $\boldsymbol{k}^{\,i,\pm}$ denotes the configuration obtained from $\boldsymbol{k}$ by replacing $k_{i}$ with $k_{i}\pm 1$, then
$w_{N}^{\pm}(i)=\frac{w}{2}\exp\!\left[-\frac{1}{2}\bigl(V_{N}(\boldsymbol{k}^{\,i,\pm})-V_{N}(\boldsymbol{k})\bigr)\right].$ (8)
Substituting Eq. (7) into this expression immediately reproduces the product form in Eq. (4). In other words, the rates are the square-root local-detailed-balance choice associated with the discrete Dyson log-gas weight $\pi_{N}^{\ast}(\boldsymbol{k})\propto e^{-V_{N}(\boldsymbol{k})}$, with the prefactor $w/2$ inherited from the bare symmetric nearest-neighbor dynamics. This also makes the reversibility with respect to $\pi_{N}^{\ast}$ transparent. As we recall below in Eq. (20), the same choice follows from the ground-state (Doob) transform of the XX chain.
The above-mentioned link with the eigenvalues $\exp{(2\pi ix_{j})}$ of random unitary matrices that perform Dyson’s Brownian motion over the unitary group $U(N)$ becomes apparent by the identification $k_{j}:=x_{j}L$ of the parameters $x_{j}$ with the rescaled particle coordinates $k_{j}$. However, we stress that here the eigenvalues are restricted to the discrete set of $L^{\rm th}$ roots of unity.
@@ 2.2 Some useful facts about the spin-1/2 XX quantum chain
$\Psi_{N}(\boldsymbol{k})=\frac{2^{N\choose 2}}{L^{N/2}}\prod_{1\leq i<j\leq N}\sin{\left(\pi\frac{k_{j}-k_{i}}{L}\right)}.$ (15)
Comparing Eq. (15) with Eq. (4), one sees directly that for a nearest-neighbor move of particle $i$,
$w_{N}^{\pm}(i)=\frac{w}{2}\,\frac{\Psi_{N}(\boldsymbol{k}^{\,i,\pm})}{\Psi_{N}(\boldsymbol{k})},$ (16)
where $\boldsymbol{k}^{\,i,\pm}$ is obtained from $\boldsymbol{k}$ by replacing $k_{i}$ with $k_{i}\pm 1$. This is precisely the off-diagonal structure generated by the ground-state transform in Eq. (20). The corresponding ground state energy is [61]
@@ 3 Hydrodynamics
Note that, in the conventional SSEP, the spectral gap of the intensity matrix scales as $O(1/L^{2})$ corresponding to diffusive scaling with dynamical exponent $z=2$. This is of course consistent with the fact that the hydrodynamic limit of the SSEP is given by the heat equation [56]. In contrast, the spectral gap of the intensity matrix of the SDEP scales as $O(1/L)$, see Eq. (18), which leads to faster relaxation with dynamical exponent $z=1$. We thus expect that, upon coarse-graining under ballistic (Eulerian) scaling, the SDEP satisfies a hydrodynamic equation resulting from the continuity equation.
@@
3.1 Heuristic derivation of the conjectured particle current (27)
We now give a heuristic derivation of the conjectured macroscopic current (27) starting from the exact microscopic expression (5). For a configuration with particle positions $k_{1},\dots,k_{N}$, let
$n_{k}^{\rm inst}=\sum_{j=1}^{N}\delta_{k,k_{j}}\in\{0,1\}$ (29)
be the instantaneous occupation variable at site $k$, and define
$S_{k}^{\rm inst}:=\sum_{q\neq k,k+1}n_{q}^{\rm inst}\log\!\left(\frac{\sin\!\bigl(\pi\frac{q-k-1}{L}\bigr)}{\sin\!\bigl(\pi\frac{q-k}{L}\bigr)}\right).$ (30)
Then Eq. (5) may be rewritten as
$j_{k}^{\rm inst}$ $=$ $\frac{w}{2}\,n_{k}^{\rm inst}(1-n_{k+1}^{\rm inst})\,e^{S_{k}^{\rm inst}}-\frac{w}{2}\,(1-n_{k}^{\rm inst})n_{k+1}^{\rm inst}\,e^{-S_{k}^{\rm inst}}.$
The key point is that $S_{k}^{\rm inst}$ contains contributions from all other particles. To separate the long-range and short-range parts, we introduce a mesoscopic cutoff $1\ll\ell\ll L$ and split the sum defining $S_{k}^{\rm inst}$ into the regions $|q-k|\leq\ell$ and $|q-k|>\ell$. For the far part, one has
$\log\!\left(\frac{\sin\!\bigl(\pi\frac{q-k-1}{L}\bigr)}{\sin\!\bigl(\pi\frac{q-k}{L}\bigr)}\right)=-\frac{\pi}{L}\cot\!\left(\pi\frac{q-k}{L}\right)+O\!\left(\frac{1}{L^{2}\sin^{2}\!\bigl(\pi\frac{q-k}{L}\bigr)}\right),$ (31)
so that, for a smooth coarse-grained density profile $\rho$,
$\sum_{|q-k|>\ell}n_{q}^{\rm inst}\log\!\left(\frac{\sin\!\bigl(\pi\frac{q-k-1}{L}\bigr)}{\sin\!\bigl(\pi\frac{q-k}{L}\bigr)}\right)=\pi\,\mathcal{H}\rho(k)+O\!\left(\frac{\ell}{L}\right)+O\!\left(\frac{1}{\ell}\right).$ (32)
Thus the contribution of particles farther than $\ell$ sites away self-averages into the non-local field $\pi\,\mathcal{H}\rho(k)$, while the remaining microscopic dependence is confined to a neighborhood of size $O(\ell)$ around the bond $(k,k+1)$. Choosing, for instance, $\ell=L^{1/2}$ makes both error terms vanish as $L\to\infty$. This suggests that, after extracting the far-field factor $e^{\pm\pi\mathcal{H}\rho}$, the remaining prefactor is local. Under local equilibrium, its coarse-grained expectation should therefore depend only on the local density $\rho(x,t)$. We are thus led to the Ansatz
@@ 3.1 Heuristic derivation of the conjectured particle current (27)
for some local functions $a_{+}$ and $a_{-}$. By reflection symmetry, $a_{+}(\rho)=a_{-}(\rho)$. Writing
$a_{+}(\rho)=a_{-}(\rho)=:\frac{1}{2}\,a(\rho),$ (34)
the current takes the form
$j[\rho](x,t)=a(\rho(x,t))\,\sinh\!\bigl(\pi\mathcal{H}\rho(x,t)\bigr).$ (35)
The same decomposition applied to the microscopic activity (obtained from (3.1) by replacing the minus sign between the two terms by a plus sign) gives the coarse-grained activity
$a[\rho](x,t)=a(\rho(x,t))\,\cosh\!\bigl(\pi\mathcal{H}\rho(x,t)\bigr).$ (36)
For a homogeneous equilibrium state of density $\rho$, one has $\mathcal{H}\rho=0$, hence
$a[\rho]=a(\rho).$ (37)
Therefore the prefactor $a(\rho)$ is naturally identified with the local equilibrium activity $a_{\rm eq}(\rho)$, which yields
$\begin{array}[]{rcl}a[\rho](x,t)&=&a_{\rm eq}(\rho(x,t))\,\cosh\!\bigl(\pi\mathcal{H}\rho(x,t)\bigr),\\[5.69054pt] j[\rho](x,t)&=&a_{\rm eq}(\rho(x,t))\,\sinh\!\bigl(\pi\mathcal{H}\rho(x,t)\bigr).\end{array}$ (38)
Finally, $a_{\rm eq}(\rho)$ is fixed by the exact lattice equilibrium result (25), which leads to the conjectured current (27).
@@ 3.2 Equivalence between non-local one-component hydrodynamics and local two-component hydrodynamics
as a local two-component hydrodynamic equation, with two densities $\rho(x)$ and $\tilde{\rho}(x)$ that satisfy a system of coupled conservation laws:
@@
The reason for this equivalence is as follows (see also the appendix of Ref. [15] for a similar calculation). The fundamental property of the Hilbert transform is that $f(x)$ and $\mathcal{H}f(x)$ are the real- and imaginary parts of the same analytic function in the upper half-plane. More specifically, for a periodic function $f(x)$ of period $L$, there exists a function $F(z)$ defined for any ${\rm Im}\,z>0$, that is analytic in the upper half-plane (including at infinity), with the same periodicity as $f$, i.e. $F(z+L)=F(z)$, and such that for all $x\in\mathbb{R}$
@@ 3.3 Relation with Abanov’s lattice free fermion imaginary-time hydrodynamics
The complex partial differential equation (43) is not new. In the context of large-scale descriptions of systems of hard-core particles or spin chains in one dimension, it was discussed by Abanov in Ref. [3]. In particular, as pointed out by Abanov, for the quantity $\zeta(z,t)=\sin P(z,t)$ , it is equivalent to a complex Hopf (or inviscid Burgers) equation:
@@
The real-time hydrodynamics of the quantum XX spin chain can be easily understood, as it reflects the dynamics of the underlying non-interacting fermions [8]. One way to proceed is to think of the non-interacting fermions semi-classically. On large scales, the state is described by a phase-space occupation function $W(x,p)$, where the momentum $p$ is defined modulo $2\pi$ because the model is considered on the lattice. Here $W$ should not be understood as a generic quantum Wigner quasi-probability, which need not be positive. Rather, in the zero-entropy semiclassical regime relevant here, $2\pi W(x,p)$ is the occupation of the phase-space cell $(x,p)$. By Pauli exclusion,
$0\leq 2\pi W(x,p)\leq 1,$ (46)
or equivalently
$0\leq W(x,p)\leq\frac{1}{2\pi}.$ (47)
The local particle density is obtained by integrating over momenta at a fixed position $x$,
@@ 4.1 Density Profiles
Here $\ln$’ stands for the principal logarithm, with the branch cut along the negative real axis, i.e. $\ln z=\ln|z|+i\,\arg z$. The two expressions above are obtained by constructing an analytic, $L$-periodic function in the upper half-plane whose boundary value satisfies ${\rm Re}\,P_{0}(x+i0^{+})=\pi\rho_{0}(x)$. For a single interval $x_{1},x_{2}]$, the ratio
$\frac{\sin\!\bigl(\frac{\pi}{L}(z-x_{1})\bigr)}{\sin\!\bigl(\frac{\pi}{L}(z-x_{2})\bigr)}$ (59)
has a phase jump of $\pi$ precisely when $x\in[x_{1},x_{2}]$, and no jump outside, so $i\log(\cdot)$ has the required real part. For two disjoint intervals, the corresponding contributions simply add, which gives the product inside the logarithm.
@@ 4.2 Limit shapes in spacetime
In Fig. 1 we see that the regions where the density is initially 0 or $1$ do not immediately melt. Instead these two regions have sharp boundaries that persist up to some time. Thus, when we look at the trajectories of particles in space-time for the SDEP, one clearly sees the appearance of a limit shape phenomenon, see Fig. 2. We observe the emergence of a sharp boundary between ‘frozen’ regions where the particle density is exactly 0 or $1$, and ‘fluctuating’ regions where the density is strictly between 0 and $1$. Such limit shape phenomena occur in various statistical physics problems, including crystal growth [66, 62], dimer coverings [42, 26, 27], statistics of random Young tableaux [57], the six-vertex model with domain-wall boundary conditions [49, 20, 63, 5], see e.g. [75, 46] for introductions. The curve separating the frozen regions and fluctuating regions is known as the arctic curve [42, 21, 19, 36, 29]. The hydrodynamic equation (39) can be exploited to make a prediction for the arctic curve for the above initial conditions, which can then be compared to the stochastic simulation of the SDEP, providing further numerical evidence for the validity of the hydrodynamic equation (39). This is what we do next.
@@
Here and below, for a polynomial $P(Z)=a_{n}Z^{n}+\cdots+a_{0}$ of degree $n$, we denote by
$\Delta(P)=(-1)^{n(n-1)/2}a_{n}^{-1}\,\mathrm{Res}(P,P^{\prime})$ (67)
its discriminant, where $\mathrm{Res}(P,P^{\prime})$ is the resultant of $P$ and its derivative $P^{\prime}$. Equivalently, $\Delta(P)=0$ if and only if $P$ and $P^{\prime}$ have a common root, i.e. if and only if $P$ has a repeated root. The resultant can be computed directly from the coefficients of $P$, for instance using the Euclidean algorithm. For a real cubic, $\Delta>0$ corresponds to three distinct real roots, whereas $\Delta<0$ corresponds to one real root and one complex-conjugate pair.